A container in a factory contains 63kL of Phosphoric acid used to produce fertilizers. How many moles of

Phosphoric acid does this container hold? The density of Phosphoric Acid is 84% higher than distilled water. 


1
Expert's answer
2021-11-02T02:07:19-0400

Solution:

Convert kL to mL:

(63 kL) × (1000 L / 1 kL) × (1000 mL / 1 L) = 63×106 mL


Density of distilled water is 1.00 g/mL

Hence,

Density of phosphoric acid = Density of distilled water + (Density of distilled water × 0.84)

Density of phosphoric acid = 1.00 g/mL + (1.00 g/mL × 0.84) = 1.84 g/mL


Density = Mass / Volume

Hence,

Mass of phosphoric acid = Density of phosphoric acid × Volume of phosphoric acid

Mass of phosphoric acid = (1.84 g/mL) × (63×106 mL) = 115.92×106 g


Moles = Mass / Molar mass

The molar mass of phosphoric acid is 97.994 g/mol

Hence,

Moles of phosphoric acid = (115.92×106 g) / (97.994 g/mol) = 1.183×106 mol

Moles of phosphoric acid = 1.18 megamol = 1.18 Mmol


Answer: 1.18 Mmol of phosphoric acid